OpenStudy (anonymous):

Factor Completely: x2 − 36 (x + 6)(x − 6) (x + 6)(x + 6) (x − 6)(x − 6) Prime

OpenStudy (anonymous):

(x+6)(x-6), because x^2 - 36 is a difference of squares. What you can do to try and figure out which one is correct is EXPAND each choice and see if you get the expression after "Factor completely:"

OpenStudy (anonymous):

\[x^2 -36\]

Parth (parthkohli):

Difference of squares. \(\Large \color{Black}{\Rightarrow a^2 - b^2 = (a + b)(a - b) \Longleftrightarrow x^2 - 6^2 = (x + 6)(x - 6) }\)

OpenStudy (anonymous):

To find out which one is correct, use the FOIL method and see which set of brackets will give you "x^2 - 36".